solve by elimination method.
A salesman sold $400 more than the rest of the sales staff. If the sales total for the day was $1600,how much did the rest of the sales staff sell?

Answer by ikleyn(52781)

.

From the condition, you have these two equations


    x - y =  400    (1)     ("A salesman sold $400 more than the rest of the sales staff")

    x + y = 1600    (2)     ("the sales total for the day was $1600")


Add the equation (it is how the Elimination method works in this case).

You will get


    2x = 2000;  hence,  x = 1000 dollars.


ANSWER.  The rest of the staff sell for 1600-1000 = 600 dollars.

Solved.